# Forecasting the Evolving Composition of Inbound Tourism Demand: A Bayesian Compositional Time Series Approach Using Platform Booking Data

### v1

@@ Abstract
Understanding how the composition of guest origin markets evolves over time is critical for destination marketing organizations, hospitality businesses, and tourism planners. We develop and apply Bayesian Dirichlet autoregressive moving average (BDARMA) models to forecast the compositional dynamics of guest origin market shares using proprietary Airbnb booking data spanning 2017–2024 across four major destination regions. Our analysis reveals substantial pandemic-induced structural breaks in origin composition, with heterogeneous recovery patterns across markets. The BDARMA framework achieves the lowest average forecast error across all destination regions, outperforming standard benchmarks including naïve forecasts, exponential smoothing, and SARIMA on log-ratio transformed data. For EMEA destinations, BDARMA achieves 23% lower forecast error than naïve methods, with statistically significant improvements. By modeling compositions directly on the simplex with a Dirichlet likelihood and incorporating seasonal variation in both mean and precision parameters, our approach produces coherent forecasts that respect the unit-sum constraint while capturing complex temporal dependencies. The methodology provides destination stakeholders with probabilistic forecasts of source market shares, enabling more informed strategic planning for marketing resource allocation, infrastructure investment, and crisis response.
@@ 1 Introduction
Tourism demand forecasting has long been recognized as essential for destination planning, resource allocation, and strategic decision-making (Song et al., 2019; Li et al., 2005; Witt and Witt, 1995). Accurate forecasts enable destination marketing organizations (DMOs) to optimize promotional spending across source markets, allow hospitality businesses to adjust pricing and staffing, and help policymakers anticipate infrastructure needs (Song and Li, 2008; Assaf et al., 2019). While the tourism forecasting literature is extensive, the vast majority of studies focus on predicting aggregate arrivals or total expenditure (Jiao and Chen, 2019; Wu et al., 2017). Considerably less attention has been paid to forecasting the composition of tourism demand—that is, how the mix of visitors from different origin markets evolves over time.
The composition of inbound tourism matters for several reasons. First, visitors from different source markets exhibit distinct spending patterns, length-of-stay distributions, and seasonal preferences (Wu et al., 2011; Divisekera, 2003). Second, the marketing channels and messaging that resonate with travelers vary substantially across cultures and geographies (Song and Li, 2008). Third, geopolitical events, exchange rate fluctuations, and public health crises affect origin markets asymmetrically, as demonstrated vividly during the COVID-19 pandemic (Gössling et al., 2020; Sigala, 2020). Understanding and forecasting compositional shifts is therefore crucial for agile destination management.
From a methodological standpoint, market share data are compositional: the shares of all origin markets must sum to unity, and each share is bounded between zero and one. Standard time series methods that ignore these constraints can produce incoherent forecasts—for example, predicted shares that sum to more than 100% or take negative values (Aitchison, 1986; Pawlowsky-Glahn et al., 2015). The compositional data analysis literature, pioneered by Aitchison (1982), offers principled approaches based on log-ratio transformations that map the simplex to unconstrained Euclidean space. However, transformed approaches can obscure interpretability and may perform poorly when shares approach zero (Greenacre, 2021).
@@
This paper contributes to both the tourism forecasting and compositional time series literatures by developing and applying BDARMA models to forecast guest origin market shares using large-scale platform booking data. We analyze Airbnb reservations from 2017 to 2024 across four major destination regions (EMEA, North America, Asia-Pacific, and Latin America), providing what is, to our knowledge, the first application of Bayesian compositional time series methods to tourism source market forecasting.
@@
We find that BDARMA models achieve the lowest average forecast error across all destination regions, substantially outperforming standard benchmarks in forecasting origin market shares. The pandemic caused dramatic compositional shifts—most notably a surge in within-region bookings at the expense of long-haul travel—with recovery trajectories that varied markedly across destination regions. Our models capture these dynamics through autoregressive and moving average terms that allow past compositions and shocks to influence current shares, while the Dirichlet likelihood ensures forecasts remain valid probability distributions. A key methodological finding is that incorporating seasonal variation in the precision parameter—not just the mean composition—substantially improves forecast accuracy.
@@ 2.1 Tourism Demand Forecasting
Tourism demand forecasting has attracted substantial scholarly attention over the past three decades (Song et al., 2019; Li et al., 2005). Early work relied primarily on econometric models—autoregressive distributed lag (ADL) specifications, error correction models, and time-varying parameter approaches—to relate arrivals or expenditure to economic determinants such as income, prices, and exchange rates (Song and Li, 2008; Witt and Witt, 1995). Song and Li (2011) and Li et al. (2006) demonstrated that incorporating cointegration and allowing for structural change improves forecast accuracy.
More recent contributions have embraced machine learning and big data. Li et al. (2020) showed that search query data from Google Trends can enhance short-term arrival forecasts. Wen et al. (2019) developed hybrid models combining ARIMA with neural networks, while Assaf et al. (2019) introduced Bayesian global vector autoregression (BGVAR) to capture spillovers across regional tourism markets. The tourism forecasting competition (Athanasopoulos et al., 2011) provided a systematic comparison of methods, finding that simple benchmarks often remain competitive with more complex approaches.
Despite this rich literature, forecasting the composition of tourism demand—as opposed to aggregate levels—has received limited attention. A notable exception is the almost ideal demand system (AIDS) approach (Wu et al., 2011; Li et al., 2006), which models budget shares allocated to different tourism products or destinations. However, AIDS models focus on expenditure allocation rather than visitor origin shares, and their linear functional form does not naturally enforce simplex constraints.
@@ 2.2 Compositional Data Analysis
Compositional data—observations that represent parts of a whole and thus sum to a constant—arise throughout the sciences (Aitchison, 1986). The standard approach transforms compositions via log-ratios (additive, centered, or isometric) to map the simplex to Euclidean space, after which conventional multivariate methods apply (Egozcue et al., 2003; Pawlowsky-Glahn et al., 2015). For time series, this enables the use of vector autoregression (VAR) and related models on transformed data (Kynčlová et al., 2015).
An alternative approach models compositions directly using distributions supported on the simplex. The Dirichlet distribution is the canonical choice, parameterized by a concentration vector $\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{C})$ with $\alpha_{c}>0$. Grunwald et al. (1993) proposed Bayesian state-space models for continuous proportions, while Zheng and Cadigan (2017) developed frequentist Dirichlet ARMA models. More recently, Katz et al. (2024) introduced the Bayesian Dirichlet ARMA (BDARMA) framework, which combines the Dirichlet likelihood with vector autoregressive moving average (VARMA) dynamics on the mean parameters. Extensions include volatility clustering via Dirichlet ARCH (Katz and Weiss, 2025) and extensive prior sensitivity analysis (Katz et al., 2025a).
The BDARMA framework offers several advantages for tourism applications. First, forecasts automatically satisfy simplex constraints without post-hoc normalization. Second, the Dirichlet concentration parameter provides a natural measure of forecast uncertainty—lower concentration implies greater dispersion across possible compositions. Third, the Bayesian approach yields full posterior predictive distributions, enabling probabilistic statements about future market shares.
@@ 2.3 Peer-to-Peer Accommodation and Platform Data
The rise of peer-to-peer (P2P) accommodation platforms, particularly Airbnb, has transformed the hospitality landscape and generated a substantial research literature (Guttentag, 2015; Dolnicar, 2019; Hall and Gössling, 2022). Studies have examined Airbnb’s competitive effects on hotels (Zervas et al., 2017), and implications for destination governance (Nieuwland and Van Melik, 2020).
Platform data offer unique advantages for tourism research. Unlike official statistics that rely on border crossings or accommodation surveys, booking data capture actual reservations with precise timing and geographic detail. Several recent studies have leveraged Airbnb data to examine pandemic-era disruptions. Katz et al. (2025b) analyzed booking lead time distributions across major U.S. cities, finding a two-phase pattern of pandemic disruption followed by incomplete recovery. Katz and Savage (2025) documented a structural shift toward longer stays, with the share of month-plus bookings nearly doubling during COVID restrictions and remaining elevated thereafter. Sainaghi and Mauri (2022) examined revenue impacts in Milan.
Our study contributes to this literature by examining a previously unexplored dimension of platform booking data: the evolving composition of guest origin markets. Understanding where guests come from—and how this mix changes over time—is fundamental for destination marketing yet has not been systematically analyzed at scale.
@@ 3.1 The BDARMA Model
where $\mathbf{X}_{t}$ is a covariate matrix (including Fourier terms for seasonality), $\boldsymbol{\beta}$ contains regression coefficients, $\mathbf{A}_{p}$ are $(C-1)\times(C-1)$ autoregressive coefficient matrices, $\mathbf{B}_{q}$ are moving average coefficient matrices, and $\tilde{\boldsymbol{\epsilon}}_{t}$ are mean-centered compositional innovations. The centering adjustment, which ensures that the MA innovations have mean zero under the Dirichlet likelihood, follows Katz (2025); technical details are provided in Appendix A.
@@ 3.2 Seasonal Precision
where $\mathbf{z}_{t}$ includes an intercept and Fourier terms ($K=6$ harmonics) capturing monthly seasonality. This specification allows compositional volatility to vary seasonally—for example, summer months may exhibit tighter concentration around expected shares due to more predictable travel patterns, while shoulder seasons may show greater dispersion. As we demonstrate in Section 5, incorporating seasonal precision substantially improves forecast accuracy compared to constant-precision specifications.
@@ 3.5 Forecasting and Evaluation
Model comparison within the BDARMA family uses leave-one-out cross-validation (LOO-CV) via Pareto-smoothed importance sampling (Vehtari et al., 2017). We report the expected log pointwise predictive density (ELPD) and effective number of parameters ($p_{\text{loo}}$). To test whether accuracy differences between methods are statistically significant, we employ the Diebold-Mariano test (Diebold and Mariano, 1995) with a heteroskedasticity and autocorrelation consistent variance estimator; details are provided in Appendix A.
@@ 4.1 Airbnb Booking Data
Our analysis uses reservation data from Airbnb’s global platform spanning January 2017 through December 2024. Airbnb is one of the world’s largest peer-to-peer accommodation marketplaces, operating in over 220 countries and regions with more than 8 million active listings (Airbnb, Inc., 2025). The platform’s standardized booking infrastructure enables consistent measurement of guest origin and destination across diverse markets.
We extract all confirmed reservations (excluding cancellations) from Airbnb’s internal bookings database. Each reservation record includes the booking date, guest country of residence, and listing location. We aggregate daily bookings to monthly frequency to smooth high-frequency noise while preserving meaningful temporal dynamics. The full sample comprises 96 months of observations.
@@ 4.2 Geographic Scope
    EMEA: Europe, Middle East, and Africa—Airbnb’s largest region by booking volume, encompassing major destinations such as France, Spain, Italy, the United Kingdom, and Germany.
@@
    NAMER: North America—comprising the United States and Canada, with the U.S. representing the platform’s founding market.
@@
    APAC: Asia-Pacific (excluding mainland China)—including Australia, Japan, South Korea, and Southeast Asian markets.
@@
    LATAM: Latin America—spanning Mexico, Brazil, Argentina, and other Central and South American countries.
@@ 4.4 Descriptive Patterns
Figure 1: Guest origin market shares by destination region, January 2017–December 2024. Stacked area charts show the compositional evolution of the top seven origin markets plus “Other” for each destination. The vertical dashed line indicates March 2020 (onset of COVID-19 pandemic). Note the dramatic compositional shifts during 2020–2021, particularly the collapse of Chinese outbound travel to APAC and the surge in within-region bookings across all destinations.
@@ 5.1 Model Selection
BDARMA(1,1) 1450.0 34.6 153.0 0.0
BDARMA(2,1) 1433.6 38.3 187.8 16.4-16.4
BDARMA(2,0) 1390.7 37.5 160.4 59.3-59.3
BDARMA(1,0) 1341.4 41.4 134.2 108.6-108.6
BDARMA(0,1) 1292.0 36.4 123.0 158.0-158.0
@@
The BDARMA(1,1) specification achieves the highest ELPD across all destination regions, indicating that both autoregressive and moving average components contribute to forecast performance. The inclusion of the MA(1) term consistently improves model fit, suggesting that compositional shocks have effects that persist beyond what the AR dynamics alone capture. Figure 5 visualizes the LOO comparison for EMEA.
Figure 5: Model comparison via leave-one-out cross-validation for EMEA. Points indicate posterior mean ELPD; error bars show $\pm 2$ standard errors. The BDARMA(1,1) specification achieves the highest expected log predictive density.
@@ 5.2 Forecast Accuracy
We evaluate forecast accuracy using rolling origin evaluation with a 3-month forecast horizon. Rolling origins begin in January 2022 and proceed in 3-month increments through October 2024, yielding 15 forecast origins with 45 total forecast observations per destination.
Table 3 presents the mean absolute error (MAE) for each method across destination regions. BDARMA achieves the lowest average MAE across all destinations (0.0072), outperforming ETS (0.0073), SARIMA (0.0079), and naïve forecasts (0.0080). For EMEA, BDARMA achieves MAE of 0.0055, representing 23% lower error than naïve forecasts (0.0072) and 6% lower than SARIMA (0.0059).
@@
BDARMA (ILR) 0.0055 0.0018 0.0122 0.0092 0.0072
SARIMA (ILR) 0.0059 0.0019 0.0154 0.0085 0.0079
ETS (ILR) 0.0066 0.0020 0.0136 0.0070 0.0073
Naïve 0.0072 0.0019 0.0124 0.0104 0.0080
Rolling Mean 0.0070 0.0028 0.0202 0.0106 0.0101
Seasonal Naïve 0.0086 0.0038 0.0326 0.0115 0.0141
Note: MAE computed as the mean absolute error averaged across components and forecast horizons. Bold indicates best performance for each destination and overall. Rolling evaluation uses origins from January 2022–October 2024 with 3-month forecast horizon.
The relative performance of methods varies systematically across destinations. BDARMA excels for EMEA and APAC, where compositional dynamics are rich—these regions feature multiple origin markets with shares between 5–25%, creating scope for autoregressive patterns to improve forecasts. For NAMER, where U.S.-origin bookings dominate at over 80%, the composition is nearly constant and all methods perform similarly. For LATAM, ETS achieves the lowest MAE, suggesting that smooth trend dynamics dominate the autoregressive patterns that BDARMA targets.
@@
Figure 6: Forecast accuracy comparison for EMEA destination. BDARMA achieves the lowest mean absolute error, followed by SARIMA and ETS. Seasonal naïve performs poorly due to pandemic-induced structural breaks that invalidate same-month-last-year patterns.
@@
$h=1$ 0.0050 0.0048 0.0051 0.0054 0.0094
$h=2$ 0.0057 0.0067 0.0067 0.0075 0.0089
$h=3$ 0.0058 0.0062 0.0080 0.0087 0.0074
Note: Bold indicates best performance at each horizon.
A notable finding is BDARMA’s strong performance in NAMER at short horizons (Table 5). Despite the region’s near-constant composition, BDARMA achieves the lowest MAE at $h=1$ (0.00098) and $h=2$ (0.00161), outperforming all benchmarks. This suggests that even in low-variation settings, the model captures subtle dynamics that simple forecasts miss.
@@
$h=1$ 0.0010 0.0013 0.0010 0.0010 0.0041
$h=2$ 0.0016 0.0018 0.0019 0.0019 0.0036
$h=3$ 0.0028 0.0026 0.0030 0.0028 0.0038
Note: Bold indicates best performance at each horizon.

### v2

# arXiv:2602.18358v2[stat.AP] 02 Mar 2026

@@ Abstract
Understanding how the composition of guest origin markets evolves over time is critical for destination marketing organizations, hospitality businesses, and tourism planners. We develop and apply Bayesian Dirichlet autoregressive moving average (BDARMA) models to forecast the compositional dynamics of guest origin market shares using proprietary Airbnb booking data spanning 2017–2025 across four major destination regions. Our analysis reveals substantial pandemic-induced structural breaks in origin composition, with heterogeneous recovery patterns across markets. In our analysis, the BDARMA framework achieves the lowest forecast error for EMEA and competitive performance across destination regions, outperforming standard benchmarks including naïve forecasts, exponential smoothing, and SARIMA on log-ratio transformed data in compositionally complex markets. For EMEA destinations, BDARMA achieves 27% lower forecast error than naïve methods ($p<0.001$), with the greatest gains where multiple origin markets compete in the 5–25% share range. By modeling compositions directly on the simplex with a Dirichlet likelihood and incorporating seasonal variation in both mean and precision parameters, our approach produces coherent forecasts that respect the unit-sum constraint while capturing complex temporal dependencies. The methodology provides destination stakeholders with probabilistic forecasts of source market shares, enabling more informed strategic planning for marketing resource allocation, infrastructure investment, and crisis response.
@@ 1 Introduction
Tourism demand forecasting has long been recognized as essential for destination planning, resource allocation, and strategic decision-making (Song et al., 2019; Li et al., 2006b; Witt and Witt, 1995). Accurate forecasts enable destination marketing organizations (DMOs) to optimize promotional spending across source markets, allow hospitality businesses to adjust pricing and staffing, and help policymakers anticipate infrastructure needs (Song and Li, 2008; Assaf et al., 2019). While the tourism forecasting literature is extensive, the vast majority of studies focus on predicting aggregate arrivals or total expenditure (Jiao and Chen, 2019; Wu et al., 2017). Considerably less attention has been paid to forecasting the composition of tourism demand, that is, how the mix of visitors from different origin markets evolves over time.
The composition of inbound tourism matters for several reasons. First, visitors from different source markets exhibit distinct spending patterns, budget allocations, and demand sensitivities (Wu et al., 2012; Divisekera, 2003). Second, the marketing channels and messaging that resonate with travelers vary substantially across cultures and geographies (Song and Li, 2008). Third, geopolitical events, exchange rate fluctuations, and public health crises affect origin markets asymmetrically, as demonstrated vividly during the COVID-19 pandemic (Gössling et al., 2021; Sigala, 2020). Understanding and forecasting compositional shifts is therefore crucial for agile destination management.
From a methodological standpoint, market share data are compositional: the shares of all origin markets must sum to unity, and each share is bounded between zero and one. Standard time series methods that ignore these constraints can produce incoherent forecasts, for example, predicted shares that sum to more than 100% or take negative values (Aitchison, 1986; Pawlowsky-Glahn et al., 2015). The compositional data analysis literature, pioneered by Aitchison (1982), offers principled approaches based on log-ratio transformations that map the simplex to unconstrained Euclidean space. However, transformed approaches can obscure interpretability and may perform poorly when shares approach zero (Greenacre, 2021).
@@
This paper contributes to both the tourism forecasting and compositional time series literatures by developing and applying BDARMA models to forecast guest origin market shares using large-scale platform booking data. We analyze Airbnb reservations from 2017 to 2025 across four major destination regions (EMEA, North America, Asia-Pacific, and Latin America), providing what is, to our knowledge, the first application of Bayesian compositional time series methods to tourism source market forecasting.
@@
We find that BDARMA models achieve the lowest forecast error for EMEA, the most compositionally complex destination region, and competitive performance elsewhere. The pandemic caused dramatic compositional shifts, most notably a surge in within-region bookings at the expense of long-haul travel, with recovery trajectories that varied markedly across destination regions. Our models capture these dynamics through autoregressive and moving average terms that allow past compositions and shocks to influence current shares, while the Dirichlet likelihood ensures forecasts remain valid probability distributions. A key methodological finding is that incorporating seasonal variation in the precision parameter, not just the mean composition, substantially improves forecast accuracy.
@@ 2.1 Tourism Demand Forecasting
Tourism demand forecasting has attracted substantial scholarly attention over the past three decades (Song et al., 2019; Li et al., 2006b). Early work relied primarily on econometric models, autoregressive distributed lag (ADL) specifications, error correction models, and time-varying parameter approaches, to relate arrivals or expenditure to economic determinants such as income, prices, and exchange rates (Song and Li, 2008; Witt and Witt, 1995). Song et al. (2011) and Li et al. (2006a) show that time-varying parameter formulations and related structural-change mechanisms can improve forecast accuracy.
More recent contributions have embraced machine learning and big data. Li et al. (2020) demonstrated that multisource big data, including search engine queries and online review platforms, can enhance short-term arrival forecasts. Wen et al. (2019) developed hybrid models that combine linear and nonlinear components (for example, ARIMA-style structure with neural networks) and leverage search query data. Assaf et al. (2019) introduced Bayesian global vector autoregression (BGVAR) to capture spillovers across regional tourism markets. The tourism forecasting competition (Athanasopoulos et al., 2011) provided a systematic comparison of methods, finding that simple benchmarks often remain competitive with more complex approaches.
Despite this rich literature, forecasting the composition of tourism demand, as opposed to aggregate levels, has received limited attention. A notable exception is the almost ideal demand system (AIDS) and related demand-system approaches (Wu et al., 2012; Li et al., 2006a), which model budget shares allocated to different tourism products, destinations, or source markets. However, demand-system approaches focus on expenditure allocation rather than visitor origin shares per se, and standard linear specifications do not naturally enforce simplex constraints.
@@ 2.2 Compositional Data Analysis
Compositional data, observations that represent parts of a whole and thus sum to a constant, arise throughout the sciences (Aitchison, 1986). The standard approach transforms compositions via log-ratios (additive, centered, or isometric) to map the simplex to Euclidean space, after which conventional multivariate methods apply (Egozcue et al., 2003; Pawlowsky-Glahn et al., 2015). For time series, this enables the use of vector autoregression (VAR) and related models on transformed data (Kynčlová et al., 2015).
An alternative approach models compositions directly using distributions supported on the simplex. The Dirichlet distribution is the canonical choice, parameterized by a concentration vector $\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{C})$ with $\alpha_{c}>0$. Grunwald et al. (1993) proposed Bayesian state-space models for continuous proportions, while Zheng and Chen (2017) developed frequentist Dirichlet ARMA models. More recently, Katz et al. (2024) introduced the Bayesian Dirichlet ARMA (BDARMA) framework, which combines the Dirichlet likelihood with vector autoregressive moving average (VARMA) dynamics on the mean parameters. Extensions include volatility clustering via Dirichlet ARCH (Katz and Weiss, 2025) and extensive prior sensitivity analysis (Katz et al., 2025a).
The BDARMA framework offers several advantages for tourism applications. First, forecasts automatically satisfy simplex constraints without post-hoc normalization. Second, the Dirichlet concentration parameter provides a natural measure of forecast uncertainty, lower concentration implies greater dispersion across possible compositions. Third, the Bayesian approach yields full posterior predictive distributions, enabling probabilistic statements about future market shares.
@@ 2.3 Peer-to-Peer Accommodation and Platform Data
The rise of peer-to-peer (P2P) accommodation platforms, particularly Airbnb, has transformed the hospitality landscape and generated a substantial research literature (Guttentag, 2015; Dolnicar, 2019; Hall et al., 2022). Studies have examined Airbnb’s competitive effects on hotels (Zervas et al., 2017), and implications for destination governance (Nieuwland and Van Melik, 2020).
Platform data offer unique advantages for tourism research. Unlike official statistics that rely on border crossings or accommodation surveys, booking data capture actual reservations with precise timing and geographic detail. Several recent studies have leveraged Airbnb data to examine pandemic-era disruptions. Katz et al. (2025b) analyzed booking lead time distributions across major U.S. cities, finding a two-phase pattern of pandemic disruption followed by incomplete recovery. Katz and Savage (2025) documented a structural shift toward longer stays, with the share of month-plus bookings nearly doubling during COVID restrictions and remaining elevated thereafter. Sainaghi and Chica-Olmo (2022) examined revenue impacts in Milan.
Our study contributes to this literature by examining a previously unexplored dimension of platform booking data: the evolving composition of guest origin markets. Understanding where guests come from, and how this mix changes over time, is fundamental for destination marketing yet has not been systematically analyzed at scale.
@@ 3.1 The BDARMA Model
where $\mathbf{X}_{t}$ is a covariate matrix (including Fourier terms for seasonality), $\boldsymbol{\beta}$ contains regression coefficients, $\mathbf{A}_{p}$ are $(C-1)\times(C-1)$ autoregressive coefficient matrices, $\mathbf{B}_{q}$ are moving average coefficient matrices, and $\tilde{\boldsymbol{\epsilon}}_{t}$ are mean-centered compositional innovations. The centering adjustment, which ensures that the MA innovations have mean zero under the Dirichlet likelihood, follows Katz (2025); technical details are provided in A.
@@ 3.2 Seasonal Precision
where $\mathbf{z}_{t}$ includes an intercept and Fourier terms ($K=6$ harmonics) capturing monthly seasonality. This specification allows compositional volatility to vary seasonally, for example, summer months may exhibit tighter concentration around expected shares due to more predictable travel patterns, while shoulder seasons may show greater dispersion. As we demonstrate in Section 5, incorporating seasonal precision substantially improves forecast accuracy compared to constant-precision specifications.
@@ 3.5 Forecasting and Evaluation
Model comparison within the BDARMA family uses leave-one-out cross-validation (LOO-CV) via Pareto-smoothed importance sampling (Vehtari et al., 2017). We report the expected log pointwise predictive density (ELPD) and effective number of parameters ($p_{\text{loo}}$). To test whether accuracy differences between methods are statistically significant, we employ the Diebold-Mariano test (Diebold and Mariano, 1995) with a heteroskedasticity and autocorrelation consistent variance estimator; details are provided in A.
@@ 4.1 Airbnb Booking Data
Our analysis uses reservation data from Airbnb’s global platform spanning January 2017 through December 2025. Airbnb is one of the world’s largest peer-to-peer accommodation marketplaces, operating in over 220 countries and regions with more than 8 million active listings (Airbnb, Inc., 2025). The platform’s standardized booking infrastructure enables consistent measurement of guest origin and destination across diverse markets.
We extract all confirmed reservations (excluding cancellations) from Airbnb’s internal bookings database. Each reservation record includes the booking date, guest country of residence, and listing location. We aggregate daily bookings to monthly frequency to smooth high-frequency noise while preserving meaningful temporal dynamics. The full sample comprises 108 months of observations.
@@ 4.2 Geographic Scope
    EMEA: Europe, Middle East, and Africa, Airbnb’s largest region by booking volume, encompassing major destinations such as France, Spain, Italy, the United Kingdom, and Germany.
@@
    NAMER: North America, comprising the United States and Canada, with the U.S. representing the platform’s founding market.
@@
    APAC: Asia-Pacific (excluding mainland China), including Australia, Japan, South Korea, and Southeast Asian markets.
@@
    LATAM: Latin America, spanning Mexico, Brazil, Argentina, and other Central and South American countries.
@@ 4.4 Descriptive Patterns
Figure 1: Guest origin market shares by destination region, January 2017–December 2025. Stacked area charts show the compositional evolution of the top seven origin markets plus “Other” for each destination. The vertical dashed line indicates March 2020 (onset of COVID-19 pandemic). Note the dramatic compositional shifts during 2020–2021, particularly the collapse of Chinese outbound travel to APAC and the surge in within-region bookings across all destinations.
@@ 5.1 Model Selection
BDARMA(2,1) 1507.1 43.1 333.3 0.0
BDARMA(1,1) 1492.2 41.1 302.1 14.9-14.9
BDARMA(2,0) 1455.7 40.9 237.3 51.4-51.4
BDARMA(1,0) 1384.0 42.3 214.0 123.1-123.1
BDARMA(0,1) 1325.0 34.4 134.6 182.1-182.1
@@
The BDARMA(1,1) specification achieves the highest ELPD for three of four destination regions (NAMER, APAC, LATAM), while BDARMA(2,1) is preferred for EMEA, indicating that both autoregressive and moving average components contribute to forecast performance. The inclusion of the MA(1) term consistently improves model fit, suggesting that compositional shocks have effects that persist beyond what the AR dynamics alone capture. Figure 5 visualizes the LOO comparison for EMEA.
Figure 5: Model comparison via leave-one-out cross-validation for EMEA. Points indicate posterior mean ELPD; error bars show $\pm 2$ standard errors. The BDARMA(2,1) specification achieves the highest expected log predictive density.
@@ 5.2 Forecast Accuracy
We evaluate forecast accuracy using rolling origin evaluation with a 6-month forecast horizon. Rolling origins begin in January 2022 and proceed in 3-month increments through April 2025, yielding 14 forecast origins with 84 total forecast observations per destination.
Table 3 presents the mean absolute error (MAE) for each method across destination regions. For EMEA, BDARMA achieves the lowest MAE (0.0060), representing 27% lower error than naïve forecasts (0.0082) and 8% lower than SARIMA (0.0065). Averaged across all four destinations, ETS achieves the lowest MAE (0.0086), followed closely by BDARMA (0.0089), SARIMA (0.0096), and naïve (0.0097).
@@
BDARMA (ILR) 0.0060 0.0023 0.0163 0.0108 0.0089
SARIMA (ILR) 0.0065 0.0022 0.0188 0.0107 0.0096
ETS (ILR) 0.0082 0.0027 0.0160 0.0075 0.0086
Naïve 0.0082 0.0025 0.0144 0.0136 0.0097
Rolling Mean 0.0074 0.0030 0.0232 0.0110 0.0112
Seasonal Naïve 0.0077 0.0035 0.0323 0.0110 0.0136
Note: MAE computed as the mean absolute error averaged across components and forecast horizons. Bold indicates best performance for each destination and overall. Rolling evaluation uses 14 origins from January 2022 through April 2025 in 3-month increments with 6-month forecast horizon ($14\times 6=84$ observations per destination).
The relative performance of methods varies systematically across destinations. BDARMA excels for EMEA, where compositional dynamics are rich, this region features multiple origin markets with shares between 5–25%, creating scope for autoregressive patterns to improve forecasts. For NAMER, where U.S.-origin bookings dominate at over 80%, the composition is nearly constant and all methods perform similarly. For LATAM, ETS achieves the lowest MAE (0.0075 vs. BDARMA’s 0.0108), suggesting that smooth trend dynamics dominate the autoregressive patterns that BDARMA targets. For APAC, naïve and ETS outperform BDARMA, likely because the post-pandemic recovery path for Chinese outbound travel is sufficiently persistent that simple extrapolation captures the dominant signal.
@@
Figure 6: Forecast accuracy comparison for EMEA destination. BDARMA achieves the lowest mean absolute error (0.0060), followed by SARIMA (0.0065), Rolling Mean (0.0074), and Seasonal Naïve (0.0077), with ETS and Naïve tied at 0.0082.
@@
$h=1$ 0.0051 0.0048 0.0053 0.0054 0.0096
$h=2$ 0.0059 0.0067 0.0069 0.0075 0.0092
$h=3$ 0.0057 0.0063 0.0082 0.0090 0.0075
Note: Bold indicates best performance at each horizon. Results shown for $h=1,2,3$; BDARMA maintains competitive accuracy at $h=4,5,6$ with similar patterns. Overall MAE across all six horizons is reported in Table 3.
For NAMER (Table 5), where U.S.-origin bookings dominate at over 80%, all methods perform similarly. At $h=1$, ETS and naïve tie for lowest MAE (0.0009); at $h=2$ ETS leads (0.0018); at $h=3$ SARIMA leads (0.0026). BDARMA remains competitive but does not dominate in this low-variation setting.
@@
$h=1$ 0.0012 0.0014 0.0009 0.0009 0.0043
$h=2$ 0.0021 0.0019 0.0018 0.0019 0.0037
$h=3$ 0.0028 0.0026 0.0029 0.0027 0.0039
Note: Bold indicates best performance at each horizon. Results shown for $h=1,2,3$; overall MAE across all six horizons is reported in Table 3.